# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ 5 Isotropic random polytopes
Next, we apply our algorithm to a common model of random convex polytopes. In Section 5.1 we first show how the expected value of the Minkowski tensors $\Phi^{0,s}_{k}(Z)$ for an isotropic random polytope $Z$ can be expressed in terms of the metric tensor $Q$ and the expected value of the intrinsic volume $V_{k}(Z)$. From this relation we then derive explicit formulas for beta-polytopes, a class of isotropic random polytopes that has proved to be useful in stochastic geometry (see, e.g., [32]). In Section 5.2, the exact evaluations of the resulting formulas are compared to the results of a simulation study based on our algorithm.
@@ 5.1 Exact expectations
(see, e.g., [27, Lemma 7] and the reference given there) we deduce that
@@
is called a beta-polytope (see [32]) in ${\mathbb{R}}^{d}$ (with parameters $\beta$ and $\ell$). Since the density $f_{d,\beta}$ is rotation invariant, the distribution of $X_{1},\ldots,X_{\ell}$ is rotation invariant, and hence $P^{\beta}_{\ell,d}$ is an isotropic random polytope. It follows from [31, Cor. 2.4] that
@@
×$\qquad\times\int_{-1}^{1}(1-h^{2})^{d\beta+\frac{d-1}{2}(d+2)}\left(\int_{-1}^{h}(1-x^{2})^{\beta+\frac{d-1}{2}}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
@@
×$\qquad\times\int_{-1}^{1}\sqrt{1-h^{2}}^{d^{2}+2d-2}\left(\int_{-1}^{h}\sqrt{1-x^{2}}^{d}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
@@ 5.2 Simulation study
Table 7: Results for 2- and 4-dimensional beta-polytopes $P_{10,d}^{\beta}$. We intersected the simulated polytope with a grid of resolution 0.005 (resp. 0.02 for $d=4$). The parameter choices were $n=50$ and $R_{n}=1$ (resp. $R_{n}=0.5$ for $d=4$). We took the average of 50 (resp. $10$ for $d=4$) renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of the expectation $\mathbb{E}\left[V_{k}\left(P_{{10},{d}}^{\beta}\right)\right]$.
$d$ $k$ $\beta$ Exact values Algorithm $d$ $k$ $\beta$ Exact values Algorithm
$2$ $1$ $\frac{1}{2}$ 2.08… 2.05(3) $4$ $3$ $\frac{1}{2}$ 0.81… 0.84(4)
$2$ $1$ $-\frac{1}{2}$ 2.57… 2.57(3) $4$ $3$ $-\frac{1}{2}$ 1.35… 1.56(8)
$2$ $2$ $\frac{1}{2}$ 1.06… 1.03(3) $4$ $4$ $\frac{1}{2}$ 0.107… 0.093(8)
$2$ $2$ $-\frac{1}{2}$ 1.71… 1.72(6) $4$ $4$ $-\frac{1}{2}$ 0.21… 0.23(2)
In this section we simulate specific beta-polytopes in dimension 2 and 4 and determine the Minkowski tensors of the resulting polytopes from the implementation of our algorithm. Since the formulas from Section 5.1 simplify for $d=2$ and $\beta\in\{-\frac{1}{2},\frac{1}{2}\}$, we focus on these two values for $\beta$ (also for $d=4$). Table 7 shows the results for the expected volume and surface area of a 2- and a 4-dimensional beta-polytope with the choice $\ell=10$. For $d=2$ we also simulated the case $\ell=100$ and obtained similar results. Furthermore, Table 8 contains the results for the expectation of the tensor $\Phi_{1}^{0,2}(P_{10,2}^{\beta})_{i,j}$ (left) resp. $\Phi_{1}^{0,4}(P_{10,2}^{\beta})_{i,j,k,l}$, and Table 9 gives the results for the expectation of the tensor $\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}$.
Table 8: Results for 2-dimensional beta-polytopes $P_{10,2}^{\beta}$. We intersected the simulated polytope with a grid of resolution 0.005. The parameter choices were $n=50$ and $R_{n}=1$. We took the average of 50 (left) resp. 10 (right) renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of
@@
$\beta$ $i$ $j$ Exact values Algorithm $\beta$ $l$ Exact values Algorithm
@@
$\frac{1}{2}$ 2 2 0.083… 0.084(2) $-\frac{1}{2}$ 1 0.0031… 0.0030(1)
$-\frac{1}{2}$ 1 1 0.102… 0.102(2) $-\frac{1}{2}$ 2 0.0031… 0.0031(1)
$-\frac{1}{2}$ 1 2 0 0.001(2)
$-\frac{1}{2}$ 2 2 0.102… 0.101(2)
@@
Table 9: Results for 4-dimensional beta-polytopes $P_{10,4}^{\beta}$. We intersected the simulated polytopes with a grid of resolution 0.02. The parameter choices were $n=50$ and $R_{n}=0.5$. We took the average of 10 renditions and computed the exact values with the formulas of Section 5.1. The table shows the values of the expectation $\mathbb{E}\left[\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}\right]$.
$\beta$ $i$ $j$ Exact values Algorithm $\beta$ $i$ $j$ Exact values Algorithm
$\frac{1}{2}$ 1 1 0.016… 0.010(2) $-\frac{1}{2}$ 1 1 0.027… 0.022(3)
$\frac{1}{2}$ 2 2 0.016… 0.011(1) $-\frac{1}{2}$ 2 2 0.027… 0.020(3)
$\frac{1}{2}$ 3 3 0.016… 0.011(1) $-\frac{1}{2}$ 3 3 0.027… 0.022(4)
$\frac{1}{2}$ 4 4 0.016… 0.014(2) $-\frac{1}{2}$ 4 4 0.027… 0.019(2)
@@ 6 Experimental data
Finally, we apply our methods to real data, more specifically, to two distinct use cases. The first example is a collection of metallic grains that can be considered to be (or at least well approximated by) sets of positive reach. The second example are nanorough surfaces that exhibit deep cuts and high ridges, i.e., they are clearly no sets of positive reach, but they can be thought of as finite unions of compact sets with positive reach. By our Theorem 3.4, we can use our methods to estimate the interfacial tensor $\Phi_{1}^{0,2}$.
6.1 Metallic grains
Figure 2: Examples of three distinct metallic grains that are represented by voxel centers. Here, each voxel center is indicated by a ball with a diameter equal to the lattice spacing, i.e., 1 $\mu$m. The underlying experimental data of the grains is from [60].
@@ 6.1 Metallic grains
Figure 3: Non-normalized histograms of the estimated volumes (a) and surface areas (b) of the metallic grains from Section 6.1. Besides a large number of small cells, we observe a few exceptionally large cells. For better visualization, only a range of values is shown, i.e., the most extreme cases are excluded here. The underlying experimental data of the grains is from [60].
@@
Figure 4: Non-normalized histograms of anisotropy indices for the metallic grains from Section 6.1. More specifically, the plots show histograms of the ratios between the smallest and largest absolute eigenvalues of the interfacial tensor $\Phi_{1}^{0,2}$ (a) and the curvature tensor $\Phi_{2}^{0,2}$ (b). The underlying experimental data of the grains is from [60].
We first apply our algorithm and code to a real data set published in [60] of a polycrystalline material, more specifically, of a nickel-based superalloy; see [60], for further details. The data set set contains, among others, 2546 individual grains, each represented by a set of voxel centers; see Figure 2, for three examples. The resolution of the grid is 1 $\mu$m. To exclude numerical artifacts, we apply a data cut and only consider grains with at least 500 voxel centers, which results in a data set with 2180 grains. For each grain, we took the average over 10 renditions. We varied $R_{n}$ between $4\cdot R_{1}$ and $24\cdot R_{1}$. The results from Figure 3 and 4 refer to the choice $R_{n}=24\cdot R_{1}$.
@@
Figure 3 shows histograms of the estimated volumes and surface areas, respectively. The distributions peak at very small values, but they exhibit comparatively strong tails. Interestingly, the tails are more pronounced for the volumes than the surface areas. We also estimated the tensors $\Phi_{k}^{0,2}$, $k\in\{0,1,2,3\}$. Figure 4 shows a histogram of the ratios between the smallest and biggest absolute eigenvalue for the tensors $\Phi_{1}^{0,2}$ and $\Phi_{2}^{0,2}$. These ratios represent scalar indices of anisotropy, which intuitively quantify the degree of anisotropy with respect to the corresponding Minkowski tensor, and which have already been successfully applied in physics [56, 57, 58, 34, 37]. For the metallic grains, these indices detect a higher degreee of anisotropy with respect to the curvature than the surface area.
@@ 6.2 Nanorough surfaces
Figure 5: Examples of nanorough surfaces from Table 10 with three different degrees of roughness, i.e., an RMS value of (a) 7 nm, (b) 24 nm, and (c) 35 nm. The data is taken from [59]; the figures have been created by Jens Uwe Neurohr.
Table 10: Estimated values of the surface and the ratio of the eigenvalues of the tensor $\Phi_{2}^{0,2}$ in comparison with triangulation results based on data from [59]. For the surface estimation with the result of Theorem 4.5 we used Remark 4.6. We used several samples with the same RMS value. The underlying experimental data of the nanorough surfaces is from [59].
Triangulation Least Squares Theorem 4.5
RMS Surface Ratio EV Surface Ratio EV Surface Ratio EV
$[\mathrm{nm}]$ $[\mu\mathrm{m}^{2}]$ $[\mu\mathrm{m}^{2}]$ $[\mu\mathrm{m}^{2}]$
7 11.6 0.285 $13.6$ 0.102 19.1 0.204
11.6 0.287 $13.5$ 0.110 19.1 0.204
11.7 0.295 $13.7$ 0.138 19.1 0.205
24 12.7 0.371 $15.6$ 0.205 19.6 0.231
12.4 0.359 $15.0$ 0.205 19.5 0.227
12.4 0.308 $15.1$ 0.098 19.6 0.229
35 12.2 0.343 $14.7$ 0.160 19.9 0.255
12.5 0.360 $15.2$ 0.184 20.0 0.259
12.3 0.326 $14.9$ 0.155 19.9 0.253
12.4 0.354 $15.0$ 0.178 19.9 0.255
@@
Next, we apply our algorithm to a distinctly different experimental data set, a nanorough surface from [59], measured via atomic force microscopy (AFM). The roughness at the nano scale, i.e., an order of magnitude above atomic resolution, was obtained by a specific etching protocol; see [59] for further details. Both protocol and three-dimensional visualizations of the data sets (see Figure 5) suggest that the nanorough surface can be described (or at least approximated) by a finite union of compact sets with positive reach.
@@
We analyze ten samples in total with three different values of the root mean square (RMS) of height distribution (7 nm, 24 nm, and 35 nm). Each sample is measured within a scan window of size $3\,\mu\text{m}\times 3\,\mu\text{m}$ and a resolution of $512\times 512$ pixels, i.e., each surface is represented by a grid of three-dimensional points, where the $z$-coordinate represents the height of the surface at the corresponding $(x,y)$ position. We varied $R_{n}$ between 12.5 and 250. The results from Table 10 refer to the value $R_{n}=25$, which is about 4 times the average nearest neighbor distance in the data. This rule of thumb balances robustness (due to a large range of radii) and accuracy (resolving fine details of the nanorough structure).
@@
We compare our estimates with those from a triangulated data set as in [59] using methods from [58]. Considering how rough our samples are, and hence challenging for our method of extrapolation, we observe a relatively good agreement of our method with the results from the triangulated data. The Voronoi estimator tends to estimate slightly larger values of the surface area and higher degrees of anisotropy (i.e., smaller values of the anisotropy index). Importantly, the trend between samples agrees for both methods.

### v2

@@ 6 Isotropic random polytopes
Computing Minkowski tensors of random polygons can be beneficial for a range of applications, examples are convex hulls in image classification and object detection [61, 80], as well as modeling and characterizing the structure of cellular materials, like polycrystalline metals, foams, or biological tissues [72, 70, 42]. Here we demonstrate the reliability of our Voronoi-LSQ algorithm also for random shapes by applying it to a prominent model of random convex polygons, known as beta-polytopes; for examples, see Fig. 8.
We recall that, in cases where it is known that the set to be estimated has positive reach (which is the case for convex sets), the Voronoi-LSQ algorithm is more suitable than the Voronoi-FD algorithm, as it requires significantly less computation time. Therefore, we restrict our attention in this section to the Voronoi-LSQ algorithm.
In Section 6.1 we first show how the expected value of the Minkowski tensors $\Phi^{0,s}_{k}(Z)$ for an isotropic random polytope $Z$ can be expressed in terms of the metric tensor $Q$ and the expected value of the intrinsic volume $V_{k}(Z)$. From this relation we then derive explicit formulas for beta-polytopes, a class of isotropic random polytopes that has proved to be useful in stochastic geometry (see, e.g., [40]). In Section 6.2, the exact evaluations of the resulting formulas are compared to the results of a simulation study based on our algorithm.
Figure 8: Three independent realizations of a two-dimensional beta-polytope $P_{10,2}^{\beta}$ for (a) $\beta=\frac{1}{2}$ and (b) $\beta=-\frac{1}{2}$.
@@ 6.1 Exact expectations
(see, e.g., [31, Lemma 7] and the reference given there) we deduce that
@@
is called a beta-polytope (see [40]) in ${\mathbb{R}}^{d}$ (with parameters $\beta$ and $\ell$). Since the density $f_{d,\beta}$ is rotation invariant, the distribution of $X_{1},\ldots,X_{\ell}$ is rotation invariant, and hence $P^{\beta}_{\ell,d}$ is an isotropic random polytope. Figure 8 shows three independent realisations of a beta-polytope $P_{10,d}^{\beta}$ for $\beta\in\{\frac{1}{2},-\frac{1}{2}\}$. It can be seen here that in the case $\beta=\frac{1}{2}$, points closer to the origin are favored compared to the case $\beta=-\frac{1}{2}$, which leads to a random polytope having a smaller expected intrinsic volume. In the following we provide explicit expressions whose numerical evaluation confirms this intuitive observation. It follows from [39, Corollary 2.4] that
@@
$\qquad\times\int_{-1}^{1}(1-h^{2})^{d\beta+\frac{d-1}{2}(d+2)}\left(\int_{-1}^{h}(1-x^{2})^{\beta+\frac{d-1}{2}}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
@@
$\qquad\times\int_{-1}^{1}\sqrt{1-h^{2}}^{d^{2}+2d-2}\left(\int_{-1}^{h}\sqrt{1-x^{2}}^{d}\,\mathrm{d}x\right)^{\ell-d}\,\mathrm{d}h.$
@@ 6.2 Simulation study
Table 7: Results of the Voronoi-LSQ algorithm for two- and four-dimensional beta-polytopes $P_{10,d}^{\beta}$. We intersected the simulated polytopes with a grid of resolution $\Delta=0.005$ for $d=2$ and $\Delta=0.02$ for $d=4$. The parameter choices were $n=50$ and $R_{n}=1$ for $d=2$ ($R_{n}=0.5$ for $d=4$). We took the average of 50 and 10 renditions for $d=2$ and $d=4$, respectively, and computed the exact values with the formulas of Section 6.1. The table shows the values of the expectation $\mathbb{E}\left[V_{k}\left(P_{{10},{d}}^{\beta}\right)\right]$.
$d$ $k$ $\beta$ Exact value Voronoi-LSQ $d$ $k$ $\beta$ Exact value Voronoi-LSQ
22 11 $\frac{1}{2}$ 2.08… 2.05(3) 44 33 $\frac{1}{2}$ 0.81… 0.84(4)
22 11 $\frac{1}{2}$ 2.57… 2.57(3) 44 33 $\frac{1}{2}$ 1.35… 1.56(8)
22 22 $\frac{1}{2}$ 1.06… 1.03(3) 44 44 $\frac{1}{2}$ 0.107… 0.093(8)
22 22 $\frac{1}{2}$ 1.71… 1.72(6) 44 44 $\frac{1}{2}$ 0.21… 0.23(2)
In this section we simulate specific beta-polytopes in dimension 2 and 4 and estimate the Minkowski tensors of the resulting polytopes by using the Voronoi-LSQ algorithm. We restrict our focus to the Voronoi-LSQ algorithm, as its applicability to convex sets, such as polytopes, is theoretically verified and it is computationally more efficient than the Voronoi-FD algorithm. Since the formulas from Section 6.1 simplify for $d=2$ and $\beta\in\{-\frac{1}{2},\frac{1}{2}\}$, we focus on these two values for $\beta$ (also for $d=4$). Table 7 shows the results for the expected volume and surface area of a two- and a four-dimensional beta-polytope with the choice $\ell=10$. For $d=2$ we also simulated the case $\ell=100$ and obtained similar results. Furthermore, Table 8 contains the results for the expectation of the tensor $\Phi_{1}^{0,2}(P_{10,2}^{\beta})_{i,j}$ (left) and $\Phi_{1}^{0,4}(P_{10,2}^{\beta})_{i,j,k,l}$ (right) respectively, and Table 9 gives the results for the expectation of the tensor $\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}$.
Table 8: Results of the Voronoi-LSQ algorithm for two-dimensional beta-polytopes $P_{10,2}^{\beta}$. We intersected the simulated polytopes with a grid of resolution $\Delta=0.005$. The parameter choices were $n=50$ and $R_{n}=1$. We took the average of 50 (left) and 10 (right) renditions and computed the exact values with the formulas of Section 6.1. The table shows the average values of $\Phi_{1}^{0,2}(P_{10,2}^{\beta})$ and $\Phi_{1}^{0,4}(P_{10,2}^{\beta})$ on the left and right side, respectively.
@@
$\beta$ $i$ $j$ Exact value Voronoi-LSQ $\beta$ $l$ Exact value Voronoi-LSQ
@@
$\frac{1}{2}$ 2 2 0.083… 0.084(2) $\frac{1}{2}$ 1 0.0031… 0.0030(1)
$\frac{1}{2}$ 1 1 0.102… 0.102(2) $\frac{1}{2}$ 2 0.0031… 0.0031(1)
$\frac{1}{2}$ 1 2 0 0.001(2)
$\frac{1}{2}$ 2 2 0.102… 0.101(2)
@@
Table 9: Results of the Voronoi-LSQ algorithm for four-dimensional beta-polytopes $P_{10,4}^{\beta}$. We intersected the simulated polytopes with a grid of resolution $\Delta=0.02$. The parameter choices were $n=50$ and $R_{n}=0.5$. We took the average of 10 renditions and computed the exact values with the formulas of Section 6.1. The table shows the values of the expectation $\mathbb{E}\left[\Phi_{3}^{0,2}(P_{10,4}^{\beta})_{i,j}\right]$.
$\beta$ $i$ $j$ Exact value Voronoi-LSQ $\beta$ $i$ $j$ Exact value Voronoi-LSQ
$\frac{1}{2}$ 1 1 0.016… 0.010(2) $\frac{1}{2}$ 1 1 0.027… 0.022(3)
$\frac{1}{2}$ 2 2 0.016… 0.011(1) $\frac{1}{2}$ 2 2 0.027… 0.020(3)
$\frac{1}{2}$ 3 3 0.016… 0.011(1) $\frac{1}{2}$ 3 3 0.027… 0.022(4)
$\frac{1}{2}$ 4 4 0.016… 0.014(2) $\frac{1}{2}$ 4 4 0.027… 0.019(2)
@@ 7 Experimental data
Finally, we apply our methods to experimental data to demonstrate their robustness in real-life scenarios. Therefore, we pick two distinct use cases, where the anisotropy of a random spatial structure can be easily related to physical (or biological) properties: a cellular structure (here represented by voxels) and a rough surface (here represented by a point cloud).
The first example consists of three-dimensional tomography data by Stinville et al. [74] for a nickel-based superalloy that forms a polycrystalline material. This material is studied under mechanical loading. Here, we are especially interested in the anisotropy of the single cells, which can be well approximated by sets of positive reach. The high resolution of this data set enables accurate estimates via our Voronoi-LSQ algorithm despite significant size disparities between cells (which can have volumes ranging from $\mu\text{m}^{3}$ to mm3).
In our second example, we challenge our method and probe it under extreme conditions; we apply it to a data set that represents black silicon, i.e., nanorough surfaces with structural features probably almost down to the atomic scale and thus even below the already high resolution of the atomic-force microscopy data from Spengler et al. [73]. The nanorough surfaces exhibit deep cuts and high ridges; hence, they can hardly be modeled by sets of positive reach (since the reach would be smaller than the resolution). Instead, these nanorough surfaces can be well approximated via the boundary of finite unions of compact sets with positive reach, namely spherical caps, because of the fabrication process via an etching procedure. We specifically want to compare our two algorithms Voronoi-FD and -LSQ for the surface area, and demonstrate consistency with previous results based on triangulations.
7.1 Metallic grains in a polycrystalline material
[…]
@@ 7.1 Metallic grains in a polycrystalline material
Figure 10: Non-normalized histograms of the estimated volumes (a) and surface areas (b) of the metallic grains from Section 7.1 by using the Voronoi-LSQ algorithm. Besides a large number of small cells, we observe a few exceptionally large cells. For better visualization, only a range of values is shown, i.e., the most extreme cases are excluded here. The underlying experimental data of the grains is from [74].
@@
Figure 11: Non-normalized histograms of anisotropy indices for the metallic grains from Section 7.1 by using the Voronoi-LSQ algorithm. More specifically, the plots show histograms of the ratios between the smallest and largest absolute eigenvalues of the interfacial tensor $\Phi_{1}^{0,2}$ (a) and the curvature tensor $\Phi_{2}^{0,2}$ (b). The underlying experimental data of the grains is from [74].
Stinville et al. [74] provides three-dimensional tomography measurements of a polycrystalline material. The data set was collected to investigate plastic deformations during mechanical loading. Such insights into structure-property relations can be used for the prediction and design of mechanical properties based on the material’s microstructure. Obviously, accurate predictions require sensitive, yet robust structural characteristics.
A natural geometric question for cellular materials (especially if subject to mechanical loading) is how to quantify anisotropy of single (deformed) cells, both the degree of anisotropy and its preferred orientations. As indicated by Fig. 1, the Minkowski tensors $\Phi_{d-1}^{0,s}$ and $\Phi_{d-2}^{0,s}$ quantify anisotropy with respect to different geometric properties, namely surface area and mean curvature, respectively.
[…]
We begin by demonstrating that our Voronoi-LSQ algorithm provides robust estimates of the volume and surface area of the cells, which vary by several orders of magnitude. Figure 10 shows histograms of the estimated volumes and surface areas, respectively. The distributions peak at very small values, but they exhibit comparatively strong tails. Interestingly, the tails are more pronounced for the volumes than the surface areas.
@@
We then quantify the anisotropy of the cells via our estimates of the tensors $\Phi_{1}^{0,2}$ and $\Phi_{2}^{0,2}$. Figure 11 shows a histogram of the ratios between the smallest and biggest absolute eigenvalues of these tensors. These ratios represent scalar indices of anisotropy, which intuitively quantify the degree of anisotropy with respect to the corresponding Minkowski tensor, and which have already been successfully applied in physics [72, 70, 71, 42, 44].
For the metallic grains, we detect via these indices a higher degree of anisotropy with respect to the curvature than the surface area. A detailed physical interpretation of these geometric findings is beyond the scope of this paper, but our results underscore the need to choose a measure of anisotropy that reflects the geometric property which is physically relevant.
@@ 7.2 Nanorough surfaces
Next, we apply our algorithms to a distinctly different experimental data set that represents a nanorough surface known as “black silicon”, from [73], measured via atomic force microscopy (AFM). The surface is represented as a point cloud that is formed by the $z$-values measured on a grid of $x$- and $y$-values.
Here, we want to estimate the surface area of the black silicon. Note that the point cloud represents the interface as a two-dimensional sheet. Therefore, at almost every point of the interface there are two normal vectors with opposite orientation, i.e., facing ‘up’ and ‘down’. Hence, $\Phi_{2}^{0,0}$ of the two-dimensional sheet is twice the value of $\Phi_{2}^{0,0}$ of the black silicon within the observation window. The latter equals half of the surface area of the black silicon within the observation window. Hence, the estimate of $\Phi_{2}^{0,0}$ of the Voronoi-LSQ algorithm can directly be used as an estimate of the surface area of the black silicon.
@@
The roughness at the nanoscale was obtained by a specific etching protocol; see [73] for further details. Both the protocol and three-dimensional visualizations of the data sets (see Figure 12) suggest that the nanorough surface can be described (or at least approximated) by a finite union of compact sets with positive reach, but it is an extreme test case for our methods because of the sharp features in the rough surface. In fact, the nanorough surface is a difficult example for any estimator of Minkowski tensors because there are features probably almost down to the atomic scale and thus below the already high resolution of about 6 nm in $x$- and $y$-direction.
Figure 12: Examples of nanorough surfaces from Figure 13 with three different degrees of roughness, i.e., an RMS value of (a) 7 nm, (b) 24 nm, and (c) 35 nm. The data is taken from [73]; the figures have been created by Jens Uwe Neurohr.
We cannot eliminate pixelation errors merely by increasing the resolution to an arbitrarily high level. Instead, the data requires accurate estimates of Minkowski tensors at a reasonable resolution. Here we want to compare our algorithms to previous results from a triangulation of the surface and show consistency between the results, as well as a robustness of our algorithm for a reasonable range of values of $R_{n}$.
Figure 13: Estimated surface areas for the different datasets as a function of $R_{n}$. The dashed lines indicate the values obtained via triangulation. The gray area corresponds to a choice of $R_{n}$ that violates the rule (4.24).
@@
Here, we analyze ten samples in total with three different values of the root mean square (RMS) of height distribution (7 nm, 24 nm, and 35 nm). Each sample is measured within a scan window of size $3\,\mu\text{m}\times 3\,\mu\text{m}$ and a resolution of $512\times 512$ pixels, i.e., each surface is represented by a grid of three-dimensional points, where the $z$-coordinate represents the height of the surface at the corresponding $(x,y)$ position.
Figure 13 shows the estimates of the surface area via the Voronoi-LSQ algorithm as a function of $R_{n}$. A proper choice of $R_{n}$ has to balance robustness (via a large range of radii) and accuracy (via a small range of radii that resolves fine details of the nanorough structure). From our test cases, we derived as a rule of thumb that $R_{n}$ should be at least four times the average nearest neighbor distance in the data; see (4.24). This rule of thumb is indicated in Fig. 13 by the gray-shaded area.
@@
For larger values of $R_{n}$, we find consistent results, i.e., no strong dependence on $R_{n}$ for a considerable range of values, roughly from $25$ to 150150\,nm. For this range, our estimated values are slightly larger than those obtained from a triangulation of the surface as in [73] using methods from [71]. Considering how rough these surfaces are and hence challenging for our method of extrapolation, we observe a relatively good agreement of our methods with the results from the triangulated data. Since the triangulation-based method tends to underestimate the surface area, our results are consistent within the accuracy of the data.
Our estimates via the Voronoi-LSQ algorithm are also consistent with those from the Voronoi-FD algorithm shown in Table 10; however, the results from the Voronoi-FD algorithm depend more strongly on the choice of the radius $\varepsilon$. If the radius is too large, a boundary correction is required for a two-dimension sheet and more importantly the parallel set hides small features of the rough surface. If the radius is too small, artifical ‘gaps’ and ‘bumps’ can lead to an overestimation of the surface area. Here, we choose a radius of twice the average nearest neighbor distance in the data. New data with an even higher resolution will be needed to compare the results of the algorithms more precisely.
Table 10: Estimated values of the surface area in $\mathrm{\mu m}^{2}$ via the Voronoi-FD algorithm in comparison with triangulation results based on data from [73]. We choose $\varepsilon=2\cdot a$ and $a$ as the average nearest neighbor distance in the data. We used several samples with the same RMS value. The underlying experimental data of the nanorough surfaces is from [73].
RMS Triangulation Voronoi-FD RMS Triangulation Voronoi-FD
7 11.6 12.6 24 12.7 13.9
11.6 12.6 12.4 13.5
11.7 12.6 12.4 13.5
35 12.2 13.4 35 12.3 13.5
12.5 13.6 12.4 13.6